OpenStudy (anonymous):

2a + 2b = c for b How do I solve for the indicated variable?

OpenStudy (anonymous):

b=c-2a over 2

OpenStudy (anonymous):

thaanks(:

OpenStudy (anonymous):

you need to isolate the variable by itself on one side. First subtract 2a from each side yielding: 2b = c - 2a then divide each side by 2: b = (c - 2a)/2

OpenStudy (anonymous):

okaay thank you .(: